in a poker hand consisting of 5 cards,find the probability of holding. (a) 3 aces, (b) 4 hearts and 1 club
Answer by Fombitz(32388)

P(A)=4/52
P(A)=3/51
P(A)=2/50
P(3A)=(4/52)(3/51)(2/50)=1/5525
.
.
.
P(H)=13/52
P(H)=12/51
P(H)=11/50
P(H)=10/49
P(C)=13/52
P(4H,C)=(13/52)(12/51)(11/50)(10/49)(13/52)=1859/2815540